OpenStudy (anonymous):

I repost question :D (old question deleted, plz give a try to this question its hard :D ) USing the mean value theorem prove that: \[2e ^{-1/4}\le \int\limits_{0}^{2}e^{-r^2-r}dr \le 2e^2\]

OpenStudy (anonymous):

mean value theorem of calculus : \[\int\limits_{a}^{b}f(x)dx = f(x)(b-a)\]

OpenStudy (zarkon):

doesn't look true to me

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

(this question is taken from ODTU's (Thecnical University of Orta Dogu ) calculus final exam ^.^

OpenStudy (zarkon):

\[2e^{-1/4}\approx1.5576\] \[\int\limits_{0}^{2}e^{-r^2-r}dr\approx.545178\]

OpenStudy (anonymous):

you see its true ^.^ but prove that by using mean value theorem :)

OpenStudy (anonymous):

but i did a mistake at mean value theorem s0zi it should be .... =f(c)(b-a)

OpenStudy (zarkon):

how is \[1.5576\le .545178\]?

OpenStudy (anonymous):

is this 0.5.... or 5.4.....??

OpenStudy (zarkon):

0.54517827464201

OpenStudy (anonymous):

ahh k now i see -.-'

OpenStudy (anonymous):

(lol @ me i red .54.... as 5.4 -.-')

OpenStudy (anonymous):

ok i give up ...

OpenStudy (zarkon):

I would give up to since\[2e ^{-1/4}\le \int\limits_{0}^{2}e^{-r^2-r}dr \le 2e^2\] is not true

OpenStudy (zarkon):

*too